A person wants to establish an annuity for retirement. He wants to make quarterly deposits for 35 years so that he can then make quarterly withdraws of $14.500.00 for 10 years. The annuity earns 7.28 % compounded quarterly.

(a) How much will have to be in the account at the time he retires?

Value of account at retirement:  
[Note: Your answer is a dollar amount and should have a dollar sign and exactly two decimal places.]

(b) How much should be deposited each quarter for 35 years in order to accumulate the required amount?

quarterly deposit:  
[Note: Your answer is a dollar amount and should have a dollar sign and exactly two decimal places.]

(c) What is the total amount of interest earned during the 45 -year period?

Total Interest Earned:  
[Note: Your answer is a dollar amount and should have a dollar sign and exactly two decimal places.]
Answer #1

a)

Quarterly withdrawal = 14500

Yearly withdrawal net = 14500*4 = 58000

PV = 58000+ 1-(140.07285-10 0.0728

= $402147.01

b)P = 402147.01 一半 0.0728 (1.07285 -1 = $2736.30

p/4 = $684.08

c) net amt paid = 2736.30 *35 = 95770.50

net amt withdrawn = 580000

interest earned = $484229.50

d)